If f(x)=5x-1, find the following:
f(a-2)

Answer by bucky(2189)

If f(x)=5x-1, find the following:
f(a-2)
Look in the sets of parentheses that follow the "f". The second set of parentheses
contains the expression a-2. The first set of parentheses just contains x. So the problem
means that there is a correspondence between x and a-2. All you do is go to the expression
.
f(x) = 5x - 1
.
and replace every x by a-2. When you do, the expression becomes:
.
f(a-2) = 5(a-2) - 1
.
Multiply out the right side so that the expression becomes:
.
f(a-2) = 5a - 10 -1
.
Finally, simplify the right side by combining the -10 and -1 to find that:
.
f(a-2) = 5a - 11
Hope this helps you to understand the basics of this problem.